A beam of light in air strikes a plate of crown glass at an angle of incidence of 40°. What is the angle of refraction?


1
Expert's answer
2021-05-24T09:00:47-0400

"n_1\sin40°=n_2\sin\beta\to"


"\beta=\sin^{-1}(n_1\sin40°/n_2)=\sin^{-1}(1\cdot\sin40°/1.52)\approx25°"